OpenStudy (anonymous):

Determine the solution of the system. y=x+6 y=2X

OpenStudy (anonymous):

x=6 Can you find Y?

OpenStudy (anonymous):

To find Y you must plug in the value of X=6 in any of the other equations

OpenStudy (anonymous):

x=6 y=12

OpenStudy (johnweldon1993):

2 equations both of which y = them you can set them equal to each other y=x+6 y=2X x + 6 = 2x -x -x x = 6 now that you know x = 6...plug that into both equations and solve for y